OpenStudy (anonymous):

Find 3√216 A. 5 B. 6 <--- C. 14 D. 15 6x6x6=216=6^3 @Blank  @Hero @IMStuck

OpenStudy (blank ):

Is the 3 small?

OpenStudy (anonymous):

@Blank  The three is small. :)

OpenStudy (anonymous):

Yeah, based on the question none of those choices seem valid

OpenStudy (anonymous):

here is an example 3*sqrt(12) is 6 sqrt(3) and here is why leave the 3 out of it. The number outside of the sqrt sign and focus on the inside. usually when you got sqrt(a) the number inside which is a can be expressed as a=b*b*c in my case 12=2*2*3 the b part which is 2 can be taken to the out side like this 3*(2)*sqrt(3) notice you only left with the c part which is 3

OpenStudy (blank ):

Answer is B. 6

OpenStudy (anonymous):

Yay! I'm right. Thanks! :)

OpenStudy (blank ):

ǝɯoɔlǝʍ ǝɹ,noʎ :)

OpenStudy (anonymous):

It is not 3 sqrt(12) but 3 sqrt(216) right?

OpenStudy (anonymous):